OpenStudy (ketz):

Graph the equation : x^2+3y^2=2...Find where it cuts the axes(x-axis and y-axis) etc...

OpenStudy (anonymous):

x^2+3y^2=2 x^2/2+3y^2/2=1 x^2/2+y^2 /2/3 =1 which is an ellipse

OpenStudy (anonymous):

x^2+3Y^2=2 it cuts x-axis where y=0 x^2=2 x=±√2 it cuts x-axis at(±√2,0) it cuts y-axis where x=0 3y^2=2 y^2=±√(3/2) it cuts y axis at(0,±√(3/2) )

OpenStudy (anonymous):